You have a sample containing 1 picomole of receptor. You also have a ligand for the receptor, which has a dissociation constant (KD) of 300 nM. If you add ligand at a concentration 300 nM, and allow the ligand and receptor to reach equilibrium, how many of the receptors are occupied?
Answer #1

Dissociation constant, Kd = 300 nm

Ligand concentration, [L] = 300 nm

Receptor concentration , [R] = 10-12 mole

Kd = [R] [L]/[R-L]

300 nm = 300 nm x 10-12 mole/[R-L]

[R-L] = 10-12 moles of receptors are occupied to the ligand = 1 picomole